For the following function find the vertical, horizontal, and oblique asymptotes, if any.
f(x)=(10x^2 - 11x - 3)/(3-5x)

Answer by lwsshak3(11628)

For the following function find the vertical, horizontal, and oblique asymptotes, if any.
f(x)=(10x^2 - 11x - 3)/(3-5x)
**
Vertical asymptotes:
set denominator=0, then solve for x
3-5x=0
5x=3
x=3/5
Vertical asymptote: x=3/5
Horizontal asymptote: none
Oblique asymptote:
Divide numerator by denominator: (by long division)
(10x^2-11x-3)/(3-5x)=-2x+1 +Remainder=6
Oblique asymptote: -2x+1 (a straight line with slope=-2 and y-intercept=1